### v1

# Conceptual Framework Toward Embodied Collective Adaptive Intelligence

Abstract
Collective Adaptive Intelligence (CAI) represent a transformative approach in artificial intelligence, wherein numerous autonomous agents collaborate, adapt, and self-organize to navigate complex, dynamic environments. This paradigm is particularly impactful in embodied AI applications, where adaptability and resilience are paramount Fan and Liu (2025). By enabling systems to reconfigure themselves in response to unforeseen challenges, CAI facilitate robust performance in real-world scenarios. This article introduces a conceptual framework for designing and analyzing CAI. It delineates key attributes including task generalization, resilience, scalability, and self-assembly, aiming to bridge theoretical foundations with practical methodologies for engineering adaptive, emergent intelligence. By providing a structured foundation for understanding and implementing CAI, this work seeks to guide researchers and practitioners in developing more resilient, scalable, and adaptable AI systems across various domains.
1 Prospectives and Challenges
Emergence and Collectivity in Complex Systems: The seminal concept of "More is Different" Anderson (1972) posits that collective behavior underpins a spectrum of phenomena spanning physics, biology, and artificial intelligence Benjamin et al. (2023). This principle—wherein complex patterns emerge from the coordinated interactions of simpler constituents—has profoundly influenced modern deep learning research Ha and Tang (2022). By abstracting collective dynamics, recent advances have sought to reconcile microscopic rules with macroscopic functionalities, such as in neural modular architectures or swarm intelligence models.
Insights from Occam’s Razor: Occam’s Razor, which prioritizes simplicity in explanatory frameworks, offers a complementary perspective on collective systems. In machine learning, models adhering to this principle often demonstrate superior generalization by mitigating overfitting to training noise Domingos (1999), Mirshekari (2023). When applied to collectivity, this heuristic suggests that systems composed of repetitive, parameter-efficient modules may outperform monolithic architectures with comparable or greater parameter counts. Such designs inherently balance expressivity and inductive bias, enabling robust performance across diverse tasks while avoiding the pitfalls of excessive algorithmic complexity.
Computational Challenges in Collective Intelligence: A critical bottleneck in scaling collective intelligent system lies in their architectural divergence from mainstream computational paradigms. While collective models may reduce absolute parameter counts, their computational demands often escalate due to recurrent interactions among components Ha and Tang (2022). This discrepancy arises from parameter interdependencies and sequential binding operations, which limit parallelization on hardware accelerators like GPUs. Consequently, optimizing collective systems necessitates rethinking not only algorithmic design but also co-adapting computational substrates to accommodate their unique operational patterns.
Collective Adaptivity: Beyond Training-Time Constraints: The concept of collective adaptivity—a paradigm emphasizing decentralized, self-organizing, self-adapting multi-agent systems—has garnered significant attention in societal systems Galesic et al. (2023) and software engineering Bucchiarone (2019), Bucchiarone and Mongiello (2019). However, its application within artificial intelligence (AI) remains relatively underexplored. This work adopts relevant definitions and posits that collective adaptive intelligence is capable of dynamically adjusting its size, topology, and functionality during inference—far surpassing the scale anticipated at training or design time. Leveraging collective adaptivity could fundamentally reshape AI architectures by decentralizing learning processes and enabling emergent, context-aware behaviors. Preliminary evidence for this capability emerges from transformative domains such as large language models, which demonstrate emergent collaboration abilities even in absense of such training Hong et al. (2024), and autonomous agents operating in open-ended environments Bauer et al. (2023). However, as of now, the emergence of collective adaptivity remains limited. Emergent behaviors that are markedly distinct from individual behaviors have yet to manifest in practice. Moreover, achieving true collective scalability requires addressing both algorithmic robustness (e.g., graceful degradation under component failure) and architectural flexibility (e.g., dynamic resource allocation), thereby bridging the gap between theoretical emergence and practical deployment.
This paper formally introduces a definition of embodied collective adaptive intelligence (CAI), delineating their defining attributes and outlining a systematic conceptual framework to contextualize their design and analysis. By establishing this foundational structure, we aim to catalyze future advancements in artificial general intelligence (AGI) research, bridging theoretical insights with scalable, interdisciplinary methodologies for engineering adaptive, emergent intelligence.
2 Collective Adaptive Intelligence: Conceptual Definitions
(a) Specialist Collective Intelligence
(b) Generalist Collective Intelligence
(c) Task and topology adaption
(d) Collective Resilience
(e) Collective Scalability
(f) Self-Assembly
Figure 1: Collective intelligences and different features of collective adaptive intelligences. Specialist: the swarm is limited to handling only one specific task and maintaining a fixed topology; Generalist: the swarm possesses the ability to instantaneously tackle various tasks and adapt to different predefined topologies; Task and topology generalization: the swarm adapts to diverse tasks and collaborative patterns; Collective resilience: the swarm reconfigures its task allocation to compensate for agent failures or loss of functionality.; Collective scalability: larger swarms enable the solution of increasingly complex tasks during testing; Self-assembly: the swarm autonomously reorganizes its collaborative structure to adapt to new tasks.
We first consider a conceptual definition of collective adaptive intelligent (CAI) system. Canonical Multi-Agent Systems (MAS) are trained on single tasks and deployed solely on the training task Weiss (1999). In traditional settings, the coordination patterns among components are predefined, which is regarded as specialist collective intelligence (SCI). However, if the MAS is pretrained on a diverse set of numerous tasks, it can potentially be deployed not only on a broad range of in-distribution tasks but also on some out-of-distribution (OOD)tasks. In such cases,it is considered as a generalist collective intelligence (GCI). Drawing inspiration from some related studies Fernandez-Marquez et al. (2013), Capodieci et al. (2016), CAI should possess at least the following features (as shown in Figure 1):
  • Task and topology adaption: Achieving generalization across diverse out-of-distribution (OOD) tasks and varying topologies, including different coordination patterns.
  • Collective Resilience: Demonstrating the ability to re-adapt and complete OOD tasks even when some components are removed from the collective.
  • Collective Scalability: Showing the capability to re-adapt and handle more complex OOD tasks when new components are added to the collective.
  • Self-Assembly: The ability to self-organize from any initial topology into an appropriate structure to accomplish OOD tasks.
Adaptation (CAI) vs. Zero-Shot Generalization (GCI): The distinction between adaptation and zero-shot generalization is frequently overlooked in discussions of model generalization. However, in partially observable Markov decision processes (POMDPs), zero-shot generalization is inherently constrained, as established in prior work on meta-learning and general-purpose systems Thrun and Pratt (1998), Kirsch et al. (2023), Fan and Liu (2025). Notably, in collective systems—where individual agents operate with limited environmental visibility—the reliance on inter-agent communication exacerbates this challenge. Here, generalization often demands adaptation to partial observability, which can directly undermine zero-shot performance. This interplay underscores a critical tension: optimizing for zero-shot capabilities may conflict with achieving robust generalization in decentralized, communication-dependent systems. Clarifying this distinction is essential for designing scalable AI frameworks in uncertain or dynamically varying environments.
3 A Conceptual Framework for Collective Adaptive Agents
We now present a conceptual framework for laying the groundwork of the foundation components for Collective Adaptive Intelligence (CAI), specifically the Collective Adaptive Agents (CAA).
We consider partial Markov decision processes $\tau=\langle\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{T},\mathcal{Z},\mathcal{R},\gamma\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $\mathcal{O}$ represents the observation space, $\mathcal{R}$ represents the utility or feedbacks, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ denotes the state transition, $\mathcal{Z}:\mathcal{S}\rightarrow\mathcal{O}$ is the observation model. The target is to maximize the accumulated utility $G(\pi)=\sum_{t}\gamma^{t}r_{t}$ through decision-making $a_{t}\sim\pi$.
We now examine the policy of a CAI system denoted as $\pi_{C}$, which comprises a set of agents $C=\{\mathcal{G}_{1},...,\mathcal{G}_{N}\}$. Each agent $\mathcal{G}_{i}$ is defined as $\langle f,h,\theta,\phi_{i,0},g_{i,0}\rangle$, where $f$ and $h$ represent predefined models and operations, while $\theta$, $\phi_{i,0}$, and $g_{i,0}$ are the parameters subject to optimization. Within this framework, each agent processes a portion of the observation $o_{i,t}$, where $o_{t}\in\mathcal{O}$ and $o_{t}=(o_{1,t},...,o_{n,t})$, and accordingly outputs a corresponding segment of the action $a_{i,t}$, with $a_{t}\in\mathcal{A}$ and $a_{t}=(a_{1,t},...,a_{n,t})$.
We now propose a conceptual framework of collective adaptive agents (CAAs) as following:
$\phi_{i,t},a_{i,t},m_{i,t},g_{i,t}=$ $f(o_{i,t},a_{i,t-1},r_{t-1},M_{i,t-1},\phi_{i,t-1},g_{i,t-1},\theta)$
$M_{i,t}=$$&\displaystyle\{(m_{j,t},g_{j,t})|\forall j,h(g_{i,t},g_{j,t})\leq h_{0}\text{ \& }$
$\text{RANK}_{j}(h(g_{i,t},g_{j,t}))<K\}$ (1)
Figure 2: A formulation of collective adaptive agents.
Equation 1 is characterized by the following features:
  • Decentralization: Equation 1 operates in a decentralized fashion.
  • Self-adapting: By observing $o_{i,t},a_{i,t-1},M_{i,t-1},r_{t-1}$ and updating $\phi_{i,t}$, each individual agent can perform self-adaptation, such as via reinforcement learning techniques, thereby adapting to specific tasks and topologies during testing.
  • Self-assembly and Collective Scalability: By utilizing $h$ and updating $g_{i,t}$, the system can reshape communication and collaboration patterns. Additionally, the number of agents can be adjusted to form systems of varying scales.
The function $f$ can be implemented using recurrent neural structures, Transformers, or linear attention mechanisms. This design allows the agent to function as an in-context learner from the outset. Additionally, $f$ can be interpreted as artificially designed optimization techniques, such as gradient descent with a specified target function. This interpretation transforms Equation 1 into a gradient-based adaptation mechanism. For the function $h$, a suitable choice might be Euclidean distance. In such a scenario, $g_{i,t}$ can be interpreted as the position (estimation) of agent $i$ in a Euclidean space. This setup draws an interesting parallel to the spatial distance between ants in a colony or between neurons in a neural system, highlighting the agent’s position and interaction dynamics within its operational space.
4 Limitations and Potential Challenges
At this stage, we propose a conceptual framework for CAI and its CAA components, though it currently lacks robust validation. Although collective intelligence has strong natural analogies and great potential for the development of artificial intelligence, mainstream computational infrastructure struggles to optimize collective intelligent systems effectively, as noted by Ha and Tang (2022). However, there is promise in exploring whether we can develop genuinely collectively scalable CAAs. The goal is to find systems that can be pre-trained at small scales or even via single-agent training, yet are capable of self-organization and cooperation at much larger scales. While prior advances in LLMs and open-world research offer some encouragement, significant further progress is still required in this area.